OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours. Age Ticket Price Under three years free Three to twelve years $32 Over twelve years $44 On a certain day, the company took 82 people on whale watching trips. There were 54 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

(82 - x)32 + (82 – 54) · 44 = C (54 + x)32 + (82 – (54 + x)) · 44 = C (82 + x)32 + (82 – (54 + x)) · 44 = C (54 – x)32 + (82 – 54) · 44 = C

OpenStudy (hoblos):

(54 – x)32 + (82 – 54) · 44 = C

OpenStudy (anonymous):

please help

OpenStudy (hoblos):

what?? do you need explanaion?

OpenStudy (ash2326):

there were total 82 people out of which 54 are children 12 years and under so there are 28 people above 12 for them the price is 44 so 44*28 x are children below 3 54-x are children between 3 and 12 so price for them=(54-x)32 C=(54-x)32+44*28